OpenStudy (anonymous):

Can someone please help me. Will reward with a follow and a medal. (precalc; i know the answer but don't know how to get there)

OpenStudy (anonymous):

1/2 - sin x =1

OpenStudy (anonymous):

the answer should be 11pi/6

OpenStudy (anonymous):

@nincompoop could you possibly help please

OpenStudy (anonymous):

@Haseeb96 could you please take a look for me

OpenStudy (butterflydreamer):

do you know the ASTC diagram? :)

OpenStudy (anonymous):

oh my im so silly of course i do :P its early morning :P

OpenStudy (haseeb96):

can u draw the question?

OpenStudy (anonymous):

I evaluated like this 1/2-sin x=1 sin x=-1/2 x=-pi/6 +2 pi n, 7pi/6 + 2 pi n

OpenStudy (haseeb96):